OpenStudy (anonymous):

what are the steps to solving square root 5 (square root 2 + square root 5)

OpenStudy (anonymous):

(sqrt(5))(sqrt(2) + sqrt(5)) distribute sqrt(5) so... sqrt(10) + 5

OpenStudy (anonymous):

that is what i got too. but the book answer is sqrt(10) + 2

OpenStudy (anonymous):

= √( 5* 2) + √5² = ....

OpenStudy (anonymous):

@peytipoo Look like you post wrong answer :(

OpenStudy (anonymous):

thats the answer the book has.

OpenStudy (anonymous):

Either your post's wrong or the book has typo! = 5 + √10

OpenStudy (anonymous):

i cant figure this one out either... √5x*√10x

OpenStudy (anonymous):

Because all the terms have same root 1/2, gather them together: = √ (5* 10 x²) = ....

OpenStudy (anonymous):

Note 50 = 2 * 25 = 2 * 5²